# Linguistic and Embedding-Based Profiling of Texts generated by Humans and Large Language Models

### v2

@@ 5.1 Text and Sentence Length
Text Length: Table 1 shows the average length of HWT and MGT in the entire dataset. Previous studies found HWT to exhibit on average longer texts than MGT (Guo et al., 2023). In our study, this tendency is confirmed with a lower margin, since some models with no repetition penalty assigned show longer texts on average than humans (e.g. gpt2). However, Section D in the Appendix presents the feature analysis across domains. Notably, in the Poetry and Book domains, HWT are longer and show greater variability than MGT, likely reflecting the effect of more “creative” domains with less rigid genre-specific constraints. This may enable human authors to express individual creativity more distinctly than models.
@@
Sentence Length: Table 1 shows the average sentence length of HWT and MGT in the entire dataset. We do not notice a clear pattern that distinguish humans and different models with decoding strategies or penalty control. Figure 1 shows that the classifier uses sentence length as a distinguishing feature, with a tendency for HWT to be longer than those of MGT.
5.2 Morphological Complexity
@@
Morphological Complexity Index (MCI): Table 1 shows the MCI for nouns and the MCI for verbs across models in the dataset. Again, we notice that human-written texts tend to be somewhat in the middle between models with different decoding strategies and the presence of penalties (e.g. mpt). However, humans rather tend to have a higher inflectional diversity for both verbs and nouns than models (See Figure 1).
5.3 Syntactic Complexity
@@
Dependency Tree Depth: Table 1 shows the average syntactical dependency depth per model across the entire dataset. We notice that humans tend to score lower in dependency depth than older models, while recent model such as gpt-4 show a more human-like amount of depth in the syntactic tree. This evolution is clear when compared to other studies in which human’s syntactic depth was lower than all the other models (Zanotto and Aroyehun, 2024).
@@
Dependency Length: To further investigate the syntactic characteristics of human-written texts and machine-generated texts, Table 1 shows the average dependency length per model across the entire dataset and complement the dependency depth results. Indeed, Humans score lower than older models like GPT-2 in dependency length and depth, while newer models produce similarly short or even shorter dependencies.
5.4 Lexical Features
@@
Word Prevalence: Table 1 shows the average word prevalence per model across the dataset. While humans score slightly lower than most models, the differences are minimal. Both humans and models exceed a two z-score threshold, indicating that 98% of the population is familiar with the words used.
@@
Type-Token Ratio: Table 1 depicts the average type-token ratio per model across the dataset. Human-written texts maintain a balanced ratio of around 0.6 indicating a balanced mix of unique words and repetitions. Models with a repetition penalty show higher type-token ratios than those (e.g. mistral-chat).
5.5 Semantic and Emotional Content
@@
Semantic Similarity: Table 1 shows the average cosine similarity per model across the dataset, measuring semantic content variation by comparing sentence-level similarity within texts. Human-written texts (HWT) exhibit a lower mean similarity than most models, indicating greater semantic diversity in human writing (See Figure 11 in Appendix). Repetition penalties help reduce redundancy, while sampling generates more diverse content than greedy decoding, as demonstrated with GPT-2.
@@
Emotionality: Table 1 shows the average emotionality score per humans and models in the entire dataset. We can notice how humans score high in emotionality in the dataset, but it does not score the highest, differently from what previous studies found (Guo et al., 2023). Indeed, difference between models seem to be driven by the decoding strategy and the presence of the penalty, where the majority of models with a sampling decoding strategy and a penalty on repetition score higher than the others (see Figure 11 in the Appendix). Looking at possible differences between emotional load for positive and negative emotions, Figure 10 in the Appendix shows how there are some exceptions where humans do not consistently score higher on negative emotions than models, contrary to previous findings (Guo et al., 2023).
@@ 5.6 Domain Variability
Figure 2 shows the variability of linguistic features across domains. The results highlight variations in the linguistic feature space, reflecting the well-established influence of genre-specific constraints on linguistic variability (Biber and Conrad, 2019). Indeed, more creative domains like poetry show more variability than less creative domains such as abstracts. Figures 6 and 7 in the Appendix show that humans exhibit more variability than machines across domains. Figure 3 shows the variability based on style embeddings across domains. Surprisingly, style embeddings capture less variability in poetry than other domains. To better understand the linguistic features potentially driving these differences, an avenue for future work will be to extract the set of individual features used to train the style embeddings model and identify which ones are responsible for the different patterns. This alternative approach is beyond the scope of the current study.
Figure 2: Domain variability with linguistic features in the entire dataset.
Figure 3: Domain variability with style embeddings in the entire dataset.
5.7 Model Variability
Figure 4 presents the variability of linguistic features across different models and their release dates. The results indicate a trend where models exhibit lower linguistic variability compared to humans and score increasingly similar to one another, suggesting a homogenization of linguistic styles of models for our set of features. Despite differences in the overall magnitude of variability, Figure 5 shows that models exhibit similar variability in style embeddings both among themselves and in comparison to humans. However, the chat models tend to exhibit variability comparable to humans and way higher than their non-chat version, as expected given the different models’ design.
Figure 4: Model Variability with linguistic features in the entire dataset across different release years.
Figure 5: Model Variability with style embeddings in the entire dataset across different release years.
@@ 6 Discussion
Our analysis of linguistic features highlights different tendencies between human-written (HWT) and machine-generated texts (MGT) in the RAID dataset. Our features importance analysis shows that human-written texts tend to have less complex syntactic structures and more varied lexical and semantic content. The tendencies of producing less complex syntactical structure is in line with previous studies (Muñoz-Ortiz et al., 2024). Our variability analysis shows that models tend to produce outputs that are very similar between themselves and are less varied than humans. We attribute this result to the importance of individual style in human-written texts. Nevertheless, models that differ in decoding strategies and repetition penalties produce outputs that reflect their intended design. Moreover, the domain variability analysis suggests an influence of linguistic constraints in less flexible and less creative domains such as abstracts (Biber and Conrad, 2019). Notably, newer models tend to be similar in term of variability between themselves for our linguistic features, possibly pointing to the phenomenon of “model collapse” (Shumailov et al., 2023), where models tend to exhibit low variance due to training on generated data. Analysis based on style embeddings reveal differing levels of variability between chat models and non-chat models, with the former exhibiting variability comparable to that of humans, while the latter show lower variability. These results can be attributed to the effect of reinforcement learning from human feedback, which is used to train chat models.
@@ 7 Conclusion and Future Work
In our analysis of linguistic features of human-written (HWT) and machine-generated texts (MGT), we show different linguistic tendencies of human-written texts and how recent models tend to generate texts with linguistic characteristics similar to humans, but exhibit a lower variability within themselves based on our set of features. Future work should explore diverse corpora with varying characteristics to verify these differences across different domains and languages. Expanding the range of linguistic features, especially those related to content such as the use of metaphors or figurative language, could provide deeper insights. Moreover, our setting of Human/Machine authorship attribution should be expanded from a binary human/non-human setting to a multi-class classification, where detection models have to attribute authorship to humans and to different LLMs.
@@ 8 Limitations
One limitation of our study lies in its generalizability. We rely on a dataset covering eight domains in the English language. Thus, the applicability of our findings is limited when considering broader linguistic variations across domains and languages. The eight domains covered in our dataset provide valuable insights, but they may not be representative of all possible linguistic contexts characteristic of all textual domains. Furthermore, texts in various languages may have unique linguistic features which limits the relevance of our results to non-English contexts.
Our analysis focuses on a set of LLMs which may already have been superseded by more advanced versions due to the rapid advancements in the field. This poses a challenge for the temporal validity of our findings as future LLMs could exhibit different linguistic patterns.
@@
Extending this study to multiple languages, datasets, and model versions could potentially enhance the applicability of our findings across broader contexts and evolving technologies. However, such an extension would require significant data curation efforts and depends on the availability of multilingual linguistic pipelines capable of characterizing texts at scale. Nevertheless, this study and our findings can serve as a foundation for further exploration of LLM outputs.
@@ 9 Ethical Considerations
Appendix A Classifiers results
This appendix section presents the results from the logistic classifier trained on a set of linguistic features.
@@ A.1 Linguistic features Classifier
Table 2: Linguistic features classifier performance across the entire dataset.
Appendix B Domain variability for humans and models in linguistic features and style detection
In this section, we show the variability within domains for humans and models considering our set of linguistic features and style embeddings. We disentangle humans and models to understand if they behave differently in terms of variability for both our linguistic features and for the style embeddings. Figure 6 show the domain variability of humans, while Figure 7 shows the domain variability of non-humans. Humans exhibit more variability than machines across domains, arguably as they are more sensible to the well-know effect of different genre-specific constraints (Biber and Conrad, 2019). However, domain variability of humans and non-humans appear way more similar when dealing with the style embeddings. Figure 8 shows how poetry is the most variable domain, arguably as the most “creative” domain of the dataset. However, even with marginal differences in terms of variability, books score last, differently from the variability of our set of linguistic features. Figure 9 confirms poetry to be the most variable domain, and exhibit marginal lower variability across several domains, for example Wikipedia or abstracts. We argue that these differences stem from the different features that are considered for characterizing the analyses. Further investigation would be needed for understanding which features trigger such different behavior.
Figure 6: Domain variability with linguistic features in the entire dataset for Humans.
Figure 7: Domain variability with linguistic features in the entire dataset for Non-humans.
Figure 8: Domain variability with style embeddings in the entire dataset for Humans.
Figure 9: Domain variability with style embeddings in the entire dataset for Non-humans.
@@ Appendix C Positive and Negative Emotions
In this section, we show the differences in positive and negative emotion loads of different models with their decoding strategy and the presence or not of a penalty on repetition. Indeed, humans score high on negative emotion load, while especially more recent models tend to cut off negative emotions, arguably due to their training processes.
@@
Figure 10: Negative and Positive Emotion Intensity per Model. This figure presents the variation in emotional intensity across humans and different models.
Appendix D Linguistic Features Visualization

### v3

@@ 5.1 How do HWT and MGT differ in terms of linguistic features?
Text Length: Table 1 shows the average length of HWT and MGT in the entire dataset. Previous studies found HWT to exhibit, on average, longer texts than MGT (Guo et al., 2023). In our study, this tendency is confirmed with a lower margin, since some models with the no repetition penalty setting generate longer texts on average than humans (e.g., gpt2). However, Section E in the Appendix presents the feature analysis across domains. Notably, in the Poetry and Book domains, HWT are longer and show greater variability than MGT, likely reflecting the effect of more “creative” domains with less rigid genre-specific constraints. This may enable human authors to express individual creativity more distinctly than models.
@@
Sentence Length: Table 1 shows the average sentence length of HWT and MGT in the entire dataset. We do not notice a clear pattern that distinguishes humans from different models with decoding strategies or penalty control. Figure 2 shows that the classifier uses sentence length as a distinguishing feature, with a tendency for HWT to be longer than those of MGT.
@@
Morphological Complexity Index (MCI): Table 1 shows the MCI for nouns and the MCI for verbs across models in the dataset. Again, we notice that human-written texts tend to be somewhat in the middle between models with different decoding strategies and the presence of penalties (e.g., mpt). However, HWT rather tend to have a higher inflectional diversity for both verbs and nouns than MGT (See Figure 2).
@@
Dependency Tree Depth: Table 1 shows the average syntactical dependency depth per model across the entire dataset. We notice that HWT tend to score lower in dependency depth than older models, while recent models such as GPT-4 show a more human-like amount of depth in the syntactic tree. This evolution is clear when compared to other studies in which the syntactic depth of HWT was lower than all the other models (Zanotto and Aroyehun, 2024).
@@
Dependency Length: To further investigate the syntactic characteristics of human-written texts and machine-generated texts, Table 1 shows the average dependency length per model across the entire dataset and complements the dependency depth results. Indeed, HWT score lower than older models like GPT-2 in dependency length and depth, while newer models produce similarly short or even shorter dependencies.
@@
Word Prevalence: Table 1 shows the average word prevalence per model across the dataset. While HWT score slightly lower than most models, the differences are minimal. Both HWT and MGT exceed a two z-score threshold, indicating that 98% of the population is familiar with the words used.
@@
Type-Token Ratio: Table 1 depicts the average type-token ratio per model across the dataset. Human-written texts maintain a balanced ratio of around 0.6, indicating a balanced mix of unique words and repetitions. Models with a repetition penalty show higher type-token ratios than those without (e.g., mistral-chat).
@@
Semantic Similarity: Table 1 shows that human-written texts (HWT) exhibit a lower mean similarity than most models, indicating greater semantic diversity in human writing (See Figure 14 in Appendix). Repetition penalties help reduce redundancy, while sampling generates more diverse content than greedy decoding, as demonstrated with GPT-2.
@@
Emotionality: In Table 1, we can notice how HWT score high in average emotionality in the dataset, but it does not score the highest, unlike what previous studies found (Guo et al., 2023). Indeed, differences between models seem to be driven by the decoding strategy and the presence of the penalty, where the majority of models with a sampling decoding strategy and a penalty on repetition score higher than the others (see Figure 14 in the Appendix). Looking at possible differences between emotional load for positive and negative emotions, Figure 10 in the Appendix shows how there are some exceptions where HWT do not consistently score higher on negative emotions than MGT, contrary to previous findings (Guo et al., 2023).
@@ 5.2 Domain Variability
Figure 3 shows the variability of linguistic features across domains. The results highlight variations in the linguistic feature space, reflecting the well-established influence of genre-specific constraints on linguistic variability (Biber and Conrad, 2019). Indeed, more creative domains like poetry show more variability than less creative domains such as abstracts. Figure 11 in Appendix shows the variability based on style embeddings across domains. Style embeddings capture less variability in poetry than in other domains. To better understand the linguistic features potentially driving these differences, an avenue for future work will be to extract the set of individual features used to train the style embeddings model and identify which ones are responsible for the different patterns. This alternative approach is beyond the scope of the current study.
Moreover, we disentangle humans and models to understand if they behave differently in terms of variability for both our linguistic features and for the style embeddings. Figure 4 shows the domain variability of HWT in the linguistic feature space, while Figure 5 shows the domain variability of MGT in the linguistic feature space. HWT exhibit more variability than MGT across domains, arguably as they are more sensible to the well-known effect of different genre-specific constraints (Biber and Conrad, 2019). Figure 4 shows how poetry is the most variable domain, arguably as the most “creative” domain of the dataset. However, the domain variability of HWT and MGT appears more similar when using the style embeddings. Figure 12 in Appendix shows how Wiki is the most variable domain, in contrast to the linguistic features analysis. However, the variation in terms of standard deviation is comparable with the results based on the linguistic features. Figure 13 in Appendix shows poetry to be the least variable domain, and exhibits marginally lower variability across several domains, for example Reddit or Abstracts. We argue that these differences stem from the different features that are considered for the analyses: explicit linguistic features vs. latent style embedding representations. Further investigation would be needed to understand which dimensions can be attributed to the different observations.
Figure 3: Domain variability with linguistic features in the entire dataset
Figure 4: Domain variability with linguistic features in the entire dataset for humans
Figure 5: Domain variability with linguistic features in the entire dataset for non-humans
Figure 6: Model Variability with linguistic features in the entire dataset across different release years
Figure 7: Model Variability with style embeddings in the entire dataset across different release years
5.3 Homogenization of Model Outputs
Figure 6 presents the variability of linguistic features across different models and their release dates. The results indicate a trend where MGT exhibit lower linguistic variability compared to HWT and score increasingly similarly to one another, suggesting a homogenization of linguistic styles of models for our set of features. Despite differences in the overall magnitude of variability, Figure 7 shows that machine-generated texts exhibit similar variability in style embeddings both among themselves and in comparison to human-written texts. However, the chat models tend to exhibit variability comparable to HWT and higher than their non-chat versions, as expected given the different design of the models.
@@ 6 Discussion
Our analysis of linguistic features highlights different tendencies between human-written (HWT) and machine-generated texts (MGT) in the RAID dataset. Our feature importance analysis shows that human-written texts tend to have less complex syntactic structures and more varied lexical and semantic content. The tendency to produce less complex syntactical structures is in line with previous studies (Muñoz-Ortiz et al., 2024). Our variability analysis shows that models tend to produce outputs that are very similar between themselves and are less varied than humans. We attribute this result to the importance of individual style in human-written texts. Nevertheless, models that differ in decoding strategies and repetition penalties produce outputs that reflect their intended design. Moreover, the domain variability analysis suggests an influence of linguistic constraints in less flexible and less creative domains such as abstracts (Biber and Conrad, 2019). Notably, newer models tend to be similar in terms of variability between themselves for our linguistic features, possibly pointing to the phenomenon of “model collapse” (Shumailov et al., 2023), where models tend to exhibit low variance due to training on generated data. Analyses based on style embeddings reveal differing levels of variability between chat models and non-chat models, with the former exhibiting variability comparable to that of humans, while the latter show lower variability. These results can be attributed to the effect of reinforcement learning from human feedback, which is used to train chat models.
@@ 7 Conclusion and Future Work
In our analysis of linguistic features of human-written (HWT) and machine-generated texts (MGT), we show different linguistic tendencies of human-written texts and how recent models tend to generate texts with linguistic characteristics similar to human-written texts, but exhibit a lower variability within themselves based on our set of features. Future work should explore diverse corpora with varying characteristics to verify these differences across different domains and languages. Expanding the range of linguistic features, especially those related to content such as the use of metaphors or figurative language, could provide deeper insights. Moreover, our setting of Human/Machine authorship attribution should be expanded from a binary human/non-human setting to a multi-class classification, where detection models have to attribute authorship to humans and to different LLMs.
@@ 8 Limitations
One limitation of our study lies in its generalizability. We rely on a dataset covering eight domains in the English language. Thus, the applicability of our findings is limited when considering broader linguistic variations across domains and languages. The eight domains covered in our dataset provide valuable insights, but they may not be representative of all possible linguistic contexts characteristic of all textual domains. Furthermore, texts in various languages may have unique linguistic features that limit the relevance of our results to non-English contexts.
While this study identifies and measures a range of linguistic features relevant for comparing human- and machine-generated texts, it does not examine how interactions between these features might provide additional insights. This represents a notable limitation, as individual features may not only operate independently but also in combination. However, analyzing such interactions poses methodological challenges due to the large number of possible feature combinations, especially in the absence of prior hypotheses or a theoretical framework indicating which interactions are likely to be salient.
Our analysis focuses on a set of LLMs that may already have been superseded by more advanced versions due to the rapid advancements in the field. This poses a challenge for the temporal validity of our findings as future LLMs could exhibit different linguistic patterns.
@@
Extending this study to multiple languages, datasets, and model versions could potentially enhance the applicability of our findings across broader contexts and evolving technologies. However, such an extension would require significant data curation efforts and depend on the availability of multilingual linguistic pipelines capable of characterizing texts at scale. Nevertheless, this study and our findings can serve as a foundation for further exploration of LLM outputs.
@@ 9 Ethical Considerations
Acknowledgments
S.E.Z. is supported by the Deutsche Forschungsgemeinschaft (DFG – German Research Foundation) under Germany’s Excellence Strategy – EXC-2035/1 – 390681379. S.A. acknowledges the support of ERC Advanced Grant 101020961 PRODEMINFO.
Appendix A Classifier results
This section presents the results from the logistic classifier and the support vector machine (SVM) trained on a set of linguistic features. We also provide a qualitative error analysis for the logistic classifier, illustrating examples of texts that were misclassified.
@@ A.1 Linguistic features Classifier
Table 2: Classification performance of the logistic regression classifier based on linguistic features
In order to demonstrate the robustness of these results, we build a second classifier using an SVM on our set of linguistic features. Table 3 shows the results from the SVM classifier in distinguishing HWT and MGT. The classifier achieves an accuracy of 0.60 across all domains, in line with the results from the logistic classifier. Moreover, Figure 8 shows the results from the feature importance analysis of the SVM model. The pattern of the feature importance analysis for the SVM classifier is very similar to that of the logistic classifier (see Figure 2 in Section 5). Particularly, all features maintain the same direction, while the coefficients are slightly lower for the SVM classifier.
Metric Non-Human Human
Accuracy 0.60
Precision 0.62 0.59
Recall 0.55 0.66
F1-score 0.58 0.62
Table 3: Classification performance of the Support Vector Machine classifier using linguistic features
Figure 8: Feature importance for the SVM classifier across all domains. A positive coefficient means that as the feature value increases, the model is more likely to predict human authorship, while a negative coefficient indicates that higher values make the model less likely to predict human authorship.
A.2 Qualitative Error Analysis
We present a qualitative error analysis of the logistic classifier by examining misclassified texts. Figure 9 displays the confusion matrix, while Table 6 provides three examples of false positives (machine-generated texts predicted as human-written texts) and false negatives (HWT predicted as MGT). To explore whether some specific features systematically drive these errors, we use Cohen’s $d$ (Cohen, 2013) to measure the standardized mean difference between misclassified and correctly classified cases. We focus on false positives (FP) versus true negatives (TN) and false negatives (FN) versus true positives (TP). This allows us to identify the features that likely drive misclassification.
Table 4 shows that false positives are primarily associated with texts that are longer and exhibit greater morphological complexity (both verbs and nouns) compared to true negatives. Thus, when machine-generated texts are long and structurally rich, the classifier tends to mistake them for HWT.
In contrast, Table 5 shows that false negatives often involve human-written texts that are shorter and morphologically simpler than true positives. Additionally, misclassified HWT tend to show higher semantic similarity, as human-written texts tend to show greater semantic diversity than MGT, according to our feature importance analysis of the logistic regression classifier.
Future work could investigate whether such cases are also challenging for human annotators, potentially leading to similar misclassification.
Figure 9: Confusion matrix showing classification performance of the logistic regresson classifier for human vs. machine-generated texts
Feature Cohen’s $d$
Text Length 1.20
Morph. Complexity (Verbs) 0.88
Morph. Complexity (Nouns) 0.81
Semantic Similarity -0.66
Emotional Load 0.52
Type–Token Ratio -0.45
Dependency Depth -0.22
Sentence Length -0.20
Dependency Length -0.18
Word Prevalence -0.16
Table 4: Cohen’s $d$ for linguistic features comparing false positives and true negatives. Positive values indicate features that are more salient in machine-generated texts (MGT) that were misclassified as human-written (HWT).
Feature Cohen’s $d$
Morph. Complexity (Verbs) -1.07
Semantic Similarity 0.92
Morph. Complexity (Nouns) -0.75
Emotional Load -0.61
Text Length -0.58
Type–Token Ratio 0.57
Dependency Length 0.46
Dependency Depth 0.44
Sentence Length 0.28
Word Prevalence 0.15
Table 5: Cohen’s $d$ for linguistic features comparing false negatives and true positives. Positive values indicate features that are more salient in HWT misclassified as MGT.
False Positives (machine-generated predicted as human-written texts)
Hey fellow Redditors, I’m on day 35 of my CT (camel trail) and I’m starting to feel a little stuck. I’ve been following the same routine for weeks now, and I can’t help but wonder if I’m clinging to my nightmare because it’s all I know at this point. […].
In this paper, we explore the connection between Brunet-Derrida particle systems, free boundary problems, and Wiener-Hopf equations. We first introduce the concept of Brunet-Derrida particle systems, which are stochastic systems that describe the evolution of a collection of particles that interact with each other through a non-local competition mechanism. […].
The trees are all so naked, and shivering with cold! They clatter over one another like drunken men when there’s wind; and the snow lies hard on their arms, as if it were trying to get them into bed—trying too hard, for they don’t move an inch in reply….. […].
False Negatives (human-written predicted as machine-generated texts)
David Innes and his captive, a member of the reptilian Mahar master race of the interior world of Pellucidar, return from the surface world in the Iron Mole invented by his friend and companion in adventure Abner Perry. Emerging in Pellucidar at an unknown location, David frees his captive. […].
The Children of Zion, published in January 1998, is considered as a documentary that was based on a collection of fragments of records compiled in Palestine in 1943 by the Eastern Center for Information, a Polish government group. […].
Newspaper columnist Mitch Albom recounts time spent with his 78-year-old sociology professor, Morrie Schwartz, at Brandeis University, who was dying from Lou Gehrig’s disease (ALS). Albom, a former student of Schwartz, had not corresponded with him since attending his college classes 16 years earlier. […].
Table 6: Examples of misclassified texts by the logistic regression classifier
Appendix B Prompt Examples
In this section, we present examples of prompts from the RAID dataset used to generate MGT. All models rely on the same specific set of prompts. For instance:
  • Write the abstract for the academic paper titled ‘‘Model Theory for a Compact Cardinal.’’
  • Write a recipe for ‘‘Anise Toasts Recipe.’’
@@ Appendix C Positive and Negative Emotions
In Figure 10, we show the differences in positive and negative emotion intensities of different models with their decoding strategy and the presence or not of a penalty on repetition. Indeed, HWT score high on negative emotion intensity, while more recent models tend to use fewer negative emotion words, arguably due to their alignment with human and/or machine feedback.
@@
Figure 10: Negative and Positive Emotion Intensity per Model. This figure presents the emotional intensity across humans and different models.
Appendix D Domain variability with style embeddings
Figure 11 shows the variability within domains for HWT and MGT based on style embeddings for the entire datasets. Moreover, Figure 12 and Figure 13 disentangle the variability respectively of HWT and MGT. We argue that differences in variability between linguistic features and style embeddings stem from the different representations that are considered for the analyses. Future research should explore potential explanations for the differing variability rankings across text domains.
Figure 11: Domain variability with style embeddings in the entire dataset
Figure 12: Domain variability with style embeddings in the entire dataset for humans
Figure 13: Domain variability with style embeddings in the entire dataset for non-humans
Appendix E Linguistic Features Visualization